For a binomial random variable with n=6 and π=0.7, P(5)=?
Answer by stanbon(75887)

For a binomial random variable with n=6 and π=0.7, P(5)=?
-----
P(x = 5) = 6C5*(0.7)^5*(0.3) = 6*0.1681*0.3 = 0.3025
====================
Cheers,
Stan H.
============